# Symmetry Analysis of the Non-Hermitian Electro-Optic Effect in Crystals

### v1

@@ V Criteria for Optical Gain
So far, we have focused solely on the linear electro-optic response of 3D materials with nontrivial Berry curvature dipoles. However, to fully understand the conditions for optical gain, it is important to also consider the material’s response in the absence of an electric bias. In particular, since these materials must exhibit a metallic response with electronic scattering Morgado et al. (2024), we should take into account the linear Drude-type response, which describes the behavior of free electrons. In this context, we assume for simplicity that the material response in the absence of bias is characterized by an isotropic Drude model. This allows us to establish a straightforward criterion for the emergence of optical gain. The permittivity of Drude’s model is given by $\varepsilon_{m}=\varepsilon_{m}^{\prime}+i\varepsilon_{m}^{\prime\prime}$, where the real $\varepsilon_{m}^{\prime}$ and imaginary $\varepsilon_{m}^{\prime\prime}$ parts are given by
$\varepsilon_{m}^{\prime}=1-\frac{\omega_{p}^{2}\tau^{2}}{\omega^{2}\tau^{2}+1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \varepsilon_{m}^{\prime\prime}=\frac{\omega_{p}^{2}\tau}{\omega^{3}\tau^{2}+\omega},$ (31)
@@
$\overline{{\bf\varepsilon}}=\varepsilon_{m}\mathbf{1}_{3\times 3}+\overline{{\bf\varepsilon}}_{\text{EO}}=\overline{{\bf\varepsilon}}^{\prime}+i\overline{{\bf\varepsilon}}^{\prime\prime}$ (32)
with $\overline{{\bf\varepsilon}}^{\prime}=\varepsilon_{m}^{\prime}\mathbf{1}_{3\times 3}+\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime}$ the Hermitian part of the response and $\overline{{\bf\varepsilon}}^{\prime\prime}=\varepsilon_{m}^{\prime\prime}\mathbf{1}_{3\times 3}+\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ the non-Hermitian part, which governs the exchange of power between the material and the wave [see Eq. (5)].
As described in previous sections, the sign of the eigenvalues of $\overline{{\bf\varepsilon}}^{\prime\prime}$ determines whether the system exhibits gain or loss. The eigenvalues of $\overline{{\bf\varepsilon}}^{\prime\prime}$ are given by $\lambda_{i}+\varepsilon_{m}^{\prime\prime}$, where $\lambda_{i}$ ($i=1,2,3$) are the eigenvalues of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$. Thus, the threshold for optical gain can be found by solving $\lambda_{i}+\varepsilon_{m}^{\prime\prime}=0$. The eigenvectors of $\overline{{\bf\varepsilon}}^{\prime\prime}$, in turn, are identical to those of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$.
@@ VI Candidate Materials
The symmetries required for the non-Hermitian linear electro-optic effect are typically found in materials such as borates Mutailipu et al. (2021), chiral perovskites Long et al. (2020), and transition metal dichalcogenides. However, for this effect to be significant, the materials must exhibit large Berry curvature dipoles and low electric conductivity Morgado et al. (2024). In this regard, Weyl semimetals (WSMs) stand out as strong candidates due to their intrinsic topological properties Ruan et al. (2016); Qian et al. (2020), which naturally give rise to large BD Facio et al. (2018); Zhang et al. (2018). Many of these WSMs fall under categories B and C making them ideal for observing the linear dichroic gain.
@@
[…]
@@ Acknowledgements.
This work was partially funded by the Institution of Engineering and Technology (IET), by the Simons Foundation (Award 733700) and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações. S.L. acknowledges FCT and IT-Coimbra for the research financial support with reference DL 57/2016/CP1353/CT000. TGR acknowledges financial support from FAPERJ, grant numbers E-26/200.959/2022 and E-26/210.100/2023, CNPq, Grant 403130/2021-2, FCT - Fundação para a Ciência e Tecnologia, project reference numbers UIDB/04650/2020, 2022.06797.PTDC and 2022.07471.CEECIND/CP1718/CT0001 (with DOI identifier: 10.54499/2022.07471.CEECIND/CP1718/ CT0001). T.A.M. acknowledges FCT for research financial support with reference CEECIND/04530/2017/CP1393/CT0004 (DOI identifier: 10.54499/CEECIND/04530/2017/CP1393/CT0004) under the CEEC Individual 2017, and IT-Coimbra for the contract as an assistant researcher with reference CT/Nº 004/2019-F00069. Work by IS was supported by Grant No. PID2021-129035NB-I00 funded by MCIN/AEI/10.13039/501100011033.
@@ Appendix B Transfer matrix for propagation along $z$
For propagation along the $z$ direction, we have $\nabla=\partial/\partial z\,\hat{\mathbf{z}}$, and these equations can be written in a matrix form as

### v2

@@ V Criteria for Optical Gain
So far, we have focused solely on the linear electro-optic response of 3D materials with nontrivial Berry curvature dipoles. However, to fully understand the conditions for optical gain, it is important to also consider the material’s response in the absence of an electric bias. In particular, since these materials must exhibit a metallic response with electronic scattering Morgado et al. (2024), we should take into account the linear Drude-type response, which describes the behavior of free electrons. In this context, we assume for simplicity that the material response in the absence of bias is characterized by an isotropic Drude model. This allows us to establish a straightforward criterion for the emergence of optical gain. The permittivity of Drude’s model is given by $\varepsilon_{\text{m}}=\varepsilon_{\text{m}}^{\prime}+i\varepsilon_{\text{m}}^{\prime\prime}$, where the real $\varepsilon_{\text{m}}^{\prime}$ and imaginary $\varepsilon_{\text{m}}^{\prime\prime}$ parts are given by
$\varepsilon_{\text{m}}^{\prime}=1-\frac{\omega_{p}^{2}\tau^{2}}{\omega^{2}\tau^{2}+1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \varepsilon_{\text{m}}^{\prime\prime}=\frac{\omega_{p}^{2}\tau}{\omega^{3}\tau^{2}+\omega},$ (31)
@@
$\overline{{\bf\varepsilon}}=\varepsilon_{\text{m}}\mathbf{1}_{3\times 3}+\overline{{\bf\varepsilon}}_{\text{EO}}=\overline{{\bf\varepsilon}}^{\prime}+i\overline{{\bf\varepsilon}}^{\prime\prime}$ (32)
with $\overline{{\bf\varepsilon}}^{\prime}=\varepsilon_{\text{m}}^{\prime}\mathbf{1}_{3\times 3}+\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime}$ the Hermitian part of the response and $\overline{{\bf\varepsilon}}^{\prime\prime}=\varepsilon_{\text{m}}^{\prime\prime}\mathbf{1}_{3\times 3}+\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ the non-Hermitian part, which governs the exchange of power between the material and the wave [see Eq. (5)].
As described in previous sections, the sign of the eigenvalues of $\overline{{\bf\varepsilon}}^{\prime\prime}$ determines whether the system exhibits gain or loss. The eigenvalues of $\overline{{\bf\varepsilon}}^{\prime\prime}$ are given by $\lambda_{i}+\varepsilon_{\text{m}}^{\prime\prime}$, where $\lambda_{i}$ ($i=1,2,3$) are the eigenvalues of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$. Thus, the threshold for optical gain can be found by solving $\lambda_{i}+\varepsilon_{\text{m}}^{\prime\prime}=0$. The eigenvectors of $\overline{{\bf\varepsilon}}^{\prime\prime}$, in turn, are identical to those of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$.
@@ VI Candidate Materials
The symmetries required for the non-Hermitian linear electro-optic effect are typically found in materials such as borates Mutailipu et al. (2021), chiral perovskites Long et al. (2020), and transition metal dichalcogenides. However, for this effect to be significant, the materials must exhibit large Berry curvature dipoles and low electric conductivity Morgado et al. (2024). In this regard, Weyl semimetals (WSMs) stand out as strong candidates due to their intrinsic topological properties Ruan et al. (2016); Qian et al. (2020), which naturally give rise to large BD Facio et al. (2018); Zhang et al. (2018). Many of these WSMs fall under categories B and C, making them ideal for observing the linear dichroic gain.
@@
In addition, materials with other symmetries are also promising candidates. For example, BiTeI, which belongs to the $3m$ symmetry group of category C, and WSM in chalcopyrites $\mathrm{CuTlSe}_{2}$, $\mathrm{AgTlTe}_{2}$, $\mathrm{AuTlTe}_{2}$, and $\mathrm{ZnPbAs}_{2}$, which belong to the $\overline{4}2m$ symmetry group and are classified in category E Ruan et al. (2016). Another example is WSMs with $\bar{4}$ symmetry Gao et al. (2021), which belong to category E. In category D, tellurium has been explored in the context of NH electro-optic responses Morgado et al. (2024).
@@ Acknowledgements.
This work was partially funded by the Institution of Engineering and Technology (IET), by the Simons Foundation (Award SFI-MPS-EWP-00008530-10) and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações. SL acknowledges FCT and IT-Coimbra for the research financial support with reference DL 57/2016/CP1353/CT000. TGR acknowledges financial support from FAPERJ, grant numbers E-26/200.959/2022 and E-26/210.100/2023, CNPq, Grant 403130/2021-2, FCT - Fundação para a Ciência e Tecnologia, project reference numbers UIDB/04650/2020, 2022.06797.PTDC and 2022.07471.CEECIND/CP1718/CT0001 (with DOI identifier: 10.54499/2022.07471.CEECIND/CP1718/ CT0001). TAM. acknowledges FCT for research financial support with reference CEECIND/04530/2017/CP1393/CT0004 (DOI identifier: 10.54499/CEECIND/04530/2017/CP1393/CT0004) under the CEEC Individual 2017, and IT-Coimbra for the contract as an assistant researcher with reference CT/Nº 004/2019-F00069. Work by IS was supported by Grant Nº PID2021-129035NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ERDF/EU.
@@ Appendix B Transfer matrix for propagation along $z$
For propagation along the $z$-direction, we have $\nabla=\partial/\partial z\,\hat{\mathbf{z}}$, and these equations can be written in a matrix form as